# Identifying and Fitting Eclipse Maps of Exoplanets with Cross-Validation

### v1

@@ 1 Introduction
The first challenge is to identify when it is justifiable to fit an eclipse map to a particular observation. In general, an eclipse map model has more degrees of freedom than a model of a planet with uniform emission. It will, therefore, normally produce a better fit to the data in an eclipse, even when the deviations from the shape of a uniform eclipse are driven by random noise, or when unresolved systematic errors produce an eclipse shape that does not depend on the true emission map. This means that it is easy to fit spurious eclipse maps to noisy or inaccurate datasets. We therefore present a statistical test using k-fold cross-validation to determine when an “eclipse mapping signal” is present. This statistical test avoids both underfitting (where the model cannot match the observations), and overfitting (where the model derives spurious information from noise).
@@
The second statistical challenge is how much complexity to allow in the fitted map, if an eclipse mapping signal is indeed present. Fitting with a high degree of spatial freedom results in highly uncertain maps with spurious small-scale features due to uncertainty and degeneracy in the observational data. Section 2 therefore presents a method to derive the appropriate spatial complexity of the fitted map, again using k-fold cross-validation. This method is based on the pixel sampling and entropy maximisation methods presented in Horne (1985) for eclipse mapping of accretion disks.
We optimise the information content of our fitted maps like Chen et al. (2024), by penalising the likelihood of the fitted model by a regularisation parameter proportional to the entropy of the map. The entropy is a function of the brightness of the pixels composing the map and corresponds to the information content of the map relative to a uniformly bright image. Welbanks et al. (2023) shows a similar method for fitting spectroscopic models to exoplanet observations using leave-one-out cross validation. Challener & Rauscher (2023) showed the application of leave-one-out cross-validation to eclipse mapping, but we chose not to use the leave-one-out method as we seek to test the predictive power of a model for a longer duration of omitted data.
We demonstrate these methods with simulated data in Section 3, showing that they correctly identify when fitting a simple low-order eclipse map is justified. We show that alternative statistical metrics are not as reliable for identifying the presence of a robust signal. We then show that, if a mapping signal is present, the method can then identify the appropriate information content for a map composed of higher-order spherical harmonics, depending on the spatial scale of the true map and the precision of the data.
@@
We then demonstrate this method for real observational data in Section 4, applying it to observations of WASP-43b with JWST MIRI/LRS, WASP-18b with JWST NIRISS/SOSS, and HD 189733b with the Spitzer Space Telescope. The method identifies a clear eclipse mapping signal for observations of WASP-43b with JWST MIRI/LRS, consistent with the findings of Hammond et al. (2024). We then derive a new eclipse map for WASP-43b, composed of higher-order spherical harmonics than the map presented in Hammond et al. (2024). Our method does not identify a significant eclipse mapping signal for observations of WASP-18b with NIRISS SOSS or the Spitzer Space Telescope observations of HD 189733b; this differs from the conclusions of Coulombe et al. (2023), Majeau et al. (2012), and de Wit et al. (2012).
@@ 2.1 Simulated time-series thermal emission
To simulate the observed thermal emission in the MIRI/LRS bandpass, the GCM simulation was post-processed using the HELIOS code (Malik et al., 2017) as described in Bell et al. (2024) to produce a map of the thermal emission averaged over the entire bandpass. We do not represent limb darkening for this latitude-longitude map, as we fit a two-dimensional map that varies only due to the rotation and occultation of the planet.
@@
To simulate the light curve with starry (Luger et al., 2019), we use the full map represented precisely by harmonics up to $\ell_{\mathrm{max}}=10$. WASP-43b is known to be tidally locked to its star (Hammond et al., 2024) so the fitted eclipse map corresponds to its permanent dayside. The simulated time series runs from orbital phase -0.6 to 0.4, covering a single orbit with an eclipse and a transit. It has 7000 data points, giving a cadence of 10.04 seconds per point, similar to the cadence of the JWST observations of WASP-43b in Bell et al. (2024).
@@ 2.2 Observed time-series thermal emission
A systematic model of instrumental effects and an astrophysical model of the orbital parameters was fitted to each dataset in each of the original studies to produce the final light curves. In this study, we do not explore the possibility of degeneracies between the systematic model, the astrophysical model, and the map model that we fit. Instead, we assume that the first two models are correctly known in advance, and only fit for the map (or a Fourier series model). In reality, there may be degeneracies between these models (Hammond et al., 2024). The JWST MIRI/LRS light curve we use for WASP-43b had its systematic and astrophysical model fitted alongside an $\ell_{\mathrm{max}}=2$ eclipse map, which should at least partially mitigate such degeneracies (Hammond et al., 2024).
@@ 2.3 Fitting thermal emission with a Fourier series
This fits the out-of-eclipse light curve as well as any eclipse mapping model. However, it does not represent the effect of non-uniform planetary flux on the eclipse ingress and egress shapes. We will use this as our null hypothesis when fitting datasets. To justify fitting an eclipse map, the map must fit the dataset better than this null hypothesis. Crucially, the map must not achieve this better fit by overfitting to noise in the eclipse shape and must actually be a better model of the true eclipse shape. We use an $\mathcal{M}(n_{\mathrm{max}}=2)$ Fourier series model as our null hypothesis as this is sufficient to capture almost all the out-of-eclipse variation of a phase curve, due to the rapid decreasing contribution to a light curve from higher-order modes (Cowan & Agol, 2008).
@@ 2.4 Fitting eclipse maps
The pixel representation and spherical harmonic representation of the map are interchangeable. Sampling the pixels makes it easier to require positive emission from the planet, as we simply place positive log-normal priors on the brightness of each pixel. These have a mean value of the peak brightness of the lightcurve divided by $\pi$, and a scale factor of 0.1 to set a very wide prior spanning multiple orders in – see the documentation of PyMC3 for the precise form of this prior (Salvatier et al., 2016) . This enforces positivity at the pixel locations only, so sometimes leads to the fitting of small areas of negative emission between pixels.
If we sampled the spherical harmonics directly instead, we could not impose positivity as a prior, and would need to measure the positivity of the map at each step while it is sampled, which we found to be computationally intensive. We also found that sampling the pixel representation detected small-scale features more effectively than the spherical harmonic representation. The dense mass matrix approach available in PyMC3 (Salvatier et al., 2016) (first implemented in Foreman-Mackey et al. (2020)) allows us to efficiently sample the degeneracies in the posterior distribution of the brightness of the pixels.
@@
For our high-order maps, designed to have high spatial freedom, we found that a spherical harmonic basis of $\ell_{\mathrm{max}}=4$ and an oversampling factor of 3 (resulting in 62 pixels to sample) produced a good compromise between spatial freedom and computational time. For our low-order maps, designed to compare directly to the null hypothesis of an $n_{\mathrm{max}}=2$ Fourier series, we use a basis of $\ell_{\mathrm{max}}=2$ and an oversampling factor of 3 (corresponding to 16 pixels). We define the pixels for a given map as a vector of brightness values $\mathbf{p}$. The sampled pixel values $\mathbf{p}$ define the vector $\mathbf{c}$ of the coefficients $c_{i}$ of the spherical harmonic representation via a matrix $\mathbf{M}$:
@@
We fit a model $\mathcal{M}$ to the data points $D$ in the light curve using PyMC3 (Salvatier et al., 2016). We sample the posterior distribution of the brightness values of the pixels $\mathbf{p}$ with 250 samples in 8 chains, thinned by a factor of 4. This ensured that the Gelman-Rubin statistic was below 1.1 for our fitted maps. Ideally we would have used more samples, but compromised with this smaller number due to the need to re-run the entire sampling process on the order of 100 times to calculate the cross-validation score. It would be possible to fit the orbital and instrumental parameters at the same time as this map (Hammond et al., 2024), but in this study we assume these parameters are known in advance.
@@ 2.7 Optimising the spatial scale of fitted maps
We include the information content of the map in the fitting process by penalising the model likelihood $p$ by a factor $2\alpha S$, following Chen et al. (2024):
@@ 3.1 Identifying eclipse mapping signals in simulated data
Figure 4 shows the eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ fitted to the $\sigma=$ 150 ppm dataset in Figure 2. The first panel shows the median of the posterior distribution of fitted maps (note that the two-dimensional structure is only constrained on the dayside). We plot the median at every point because we sample the pixels composing the map; this plot therefore shows the median value of the sampled posterior distribution. The next two panels of Figure 4 show slices of the posterior distribution of the fitted maps on the dayside through the substellar point, as functions of longitude and latitude. The true map is composed of many high-order spherical harmonic components beyond $\ell=2$, so this fitted $\ell_{\mathrm{max}}=2$ map cannot match the shape of the true map very closely. It does, at least, fit the residual signal accurately, so we suggest that this map is statistically supported, if not especially accurate.
@@
The first row of Figure 3 shows the simulated data for $\sigma=$ 150 ppm, with the Fourier series and eclipse map models fitted using an incorrect value for the transit time $t_{0}^{\prime}=t_{0}+10\mathrm{s}$. As we assume a circular orbit, this produces the same offset of $+10\mathrm{s}$ in the assumed eclipse time. This simple change produces a systematic error in the astrophysical model, which means that the eclipse mapping signal and the eclipse map are fitted poorly regardless of the precision of the data. An error in assumed eclipse time could also arise in other ways, such as a slightly eccentric orbit producing an offset in eclipse timing relative to a measured transit time. Fitting the model with this delayed eclipse timing results in an eclipse ingress that is too high, and an eclipse egress that is too low. The model wrongly adjusts the eclipse map to compensate for this error.
@@ 3.3 Optimising the information content of fitted maps
Having used a simple eclipse mapping model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ to establish the presence of an eclipse mapping signal in the dataset with $\sigma=$ 150 ppm in Figure 2, we can now fit a map with more complex structure. Figure 7 shows an eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=4)$ fitted to this dataset using spherical harmonics up to order $\ell_{\mathrm{max}}=4$. This high-order map is able to fit the true map in theory, but in practice it overfits a wide range of solutions, resulting in an impractically large uncertainty.
@@
Figure 9: The thermal emission from 5 to 10.5 $\mu$m from the hot Jupiter WASP-43b observed by JWST MIRI/LRS (Bell et al., 2024). The layout of panels is the same as Figure 2. The black points show the first eclipse and the grey points show the second eclipse, binned every 10 points. The eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ fits the eclipse shape much better than the Fourier series model $\mathcal{M}(n_{\mathrm{max}}=2)$. This is confirmed by the higher cross-validation score for the eclipse map model for this dataset shown in Figure 12 and listed in Table 2. The data points in the right-hand panel appear slightly different to the data points in Hammond et al. (2024) because here we bin them with a slightly longer time cadence.

### v2

@@ 1 Introduction
The first challenge is to identify when it is justifiable to fit an eclipse map to a particular observation. In general, an eclipse map model has more degrees of freedom than a model of a planet with uniform emission. It will, therefore, normally produce a better fit to the data in an eclipse, even when the deviations from the shape of a uniform eclipse are driven by random noise, or when unresolved systematic errors produce an eclipse shape that does not depend on the true emission map. This means that it is easy to fit spurious eclipse maps to noisy or inaccurate datasets. Section 2 therefore presents a statistical test using k-fold cross-validation to determine when an “eclipse mapping signal” is present. This statistical test avoids both underfitting (where the model cannot match the observations), and overfitting (where the model derives spurious information from noise). Welbanks et al. (2023) shows a similar method for fitting spectroscopic models to exoplanet observations using leave-one-out cross validation. Challener et al. (2023) showed the application of leave-one-out cross-validation to eclipse mapping, but we chose not to use the leave-one-out method as we seek to test the predictive power of a model for a longer duration of omitted data.
@@
The second statistical challenge is how much complexity to allow in the fitted map, if an eclipse mapping signal is indeed present. Fitting with a high degree of spatial freedom results in highly uncertain maps with spurious small-scale features due to uncertainty and degeneracy in the observational data. de Wit et al. (2012) investigated this issue issue, showing how fitting an eclipse map with a model with more small-scale freedom (a single hot-spot of variable size) resulted in greater uncertainty than a model with only large-scale freedom (a low-order spherical harmonic map). Rauscher et al. (2018) developed the method of “eigenmapping” to address this issue, orthogonalising the light-curves used to fit an observation, and then determining the number of corresponding maps used to fit the data by optimising the Bayesian Information Criterion. In Section 2 we present a new eclipse mapping method to derive the appropriate spatial complexity of the fitted map, again using k-fold cross-validation. This method is based on the pixel sampling and entropy maximisation methods presented in Horne (1985) and Chen et al. (2024) for mapping of accretion disks and brown dwarfs.
@@
We demonstrate these methods with simulated data in Section 3, showing that they correctly identify when fitting a simple low-order eclipse map is justified. We show that the method can then identify the appropriate information content for a map composed of higher-order spherical harmonics, depending on the spatial scale of the true map and the precision of the data. We then demonstrate this method for real observational data in Section 4, applying it to observations of WASP-43b with JWST MIRI/LRS, WASP-18b with JWST NIRISS/SOSS, and HD 189733b with the Spitzer Space Telescope. The method identifies a clear eclipse mapping signal for observations of WASP-43b with JWST MIRI/LRS, consistent with the findings of Hammond et al. (2024), and fits a new map optimised by our new method. Our method does not identify a significant eclipse mapping signal for observations of WASP-18b with NIRISS SOSS or the Spitzer Space Telescope observations of HD 189733b; this differs from the conclusions of Coulombe et al. (2023), Majeau et al. (2012), and de Wit et al. (2012).
@@ 2.1 Simulated time-series thermal emission
To simulate the observed thermal emission in the MIRI/LRS bandpass, the GCM simulation was post-processed using the HELIOS code (Malik et al., 2017) to produce a map of the planetary thermal flux averaged over the entire bandpass, normalised by a stellar spectrum from the PHOENIX model database (Allard & Hauschildt, 1995; Husser et al., 2013) as described in Mendonça et al. (2018b). We do not represent limb darkening for this latitude-longitude map, as we fit a two-dimensional map that varies only due to the rotation and occultation of the planet.
@@
To simulate the light curve with starry (Luger et al., 2019), we use the full map represented precisely by harmonics up to $\ell_{\mathrm{max}}=10$. WASP-43b is assumed to be tidally locked to its star, based on its short orbital period and consistent thermal phase curve amplitude, (Hellier et al., 2011) so the fitted eclipse map corresponds to its permanent dayside. The simulated time series runs from orbital phase -0.6 to 0.4, covering a single orbit with an eclipse and a transit. It has 7000 data points, giving a cadence of 10.04 seconds per point, similar to the cadence of the JWST observations of WASP-43b in Bell et al. (2024).
@@ 2.2 Observed time-series thermal emission
A systematic model of instrumental effects and an astrophysical model of the system parameters was fitted to each dataset in each of the original studies to produce the final light curves (Majeau et al., 2012; Coulombe et al., 2023; Hammond et al., 2024). In this study, we do not explore the possibility of degeneracies between the systematic model, the astrophysical model, and the map model that we fit. Instead, we assume that the first two models are correctly known in advance, and only fit for the map (or a Fourier series model). In reality, there may be degeneracies between these models (de Wit et al., 2012; Hammond et al., 2024). The JWST MIRI/LRS light curve we use for WASP-43b had its systematic and astrophysical model fitted alongside an $\ell_{\mathrm{max}}=2$ eclipse map, which should at least partially mitigate such degeneracies (Hammond et al., 2024).
@@ 2.3 Fitting thermal emission with a Fourier series
This fits the out-of-eclipse light curve as well as any eclipse mapping model. However, it does not represent the effect of non-uniform planetary flux on the eclipse ingress and egress shapes. Strictly speaking, the eclipse shape in this model is slightly modified from the eclipse shape of a uniform planet by the small change in the Fourier series model over the short duration of the eclipse.
We will use this Fourier series as our null hypothesis when fitting datasets. To justify fitting an eclipse map, the map must fit the dataset better than this null hypothesis. Crucially, the map must not achieve this better fit by overfitting to noise in the eclipse shape and must actually be a better model of the true eclipse shape. We use an $\mathcal{M}(n_{\mathrm{max}}=2)$ Fourier series model as our null hypothesis as this is sufficient to capture almost all the out-of-eclipse variation of a phase curve, due to the rapidly decreasing contribution to a light curve from higher-order modes (Cowan & Agol, 2008).
@@ 2.4 Fitting eclipse maps
The pixel representation and spherical harmonic representation of the map are interchangeable. Sampling the pixels makes it easier to require positive emission from the planet, as we simply place positive log-normal priors on the brightness of each pixel. These have a mean value of the peak brightness of the lightcurve divided by $\pi$, and a scale factor of 0.1 to set a very wide prior spanning multiple orders of magnitude in planetary flux (see the documentation of PyMC3 for the precise form of this prior (Salvatier et al., 2016)). This enforces positivity at the pixel locations only, so sometimes leads to the fitting of small areas of negative emission between pixels.
If we sampled the spherical harmonics directly instead, we could not impose positivity as a prior, and would need to measure the positivity of the map at each step while it is sampled, which we found to be computationally intensive. We also found qualitatively that sampling pixels detected small-scale features more effectively than sampling the spherical harmonics. We suggest that this could be because each parameter in the pixel representation corresponds directly to a spatial region, so a small localised feature can be added by varying one parameter only (or a small number of parameters). On the other hand, introducing a new small-scale feature when sampling the spherical harmonic basis may require the sampler to adjust the values of all the spherical harmonic coefficients, as each one affects the entire map. The dense mass matrix approach available in PyMC3 (Salvatier et al., 2016) (first implemented in Foreman-Mackey et al. (2020)) allows us to efficiently sample the degeneracies in the posterior distribution of the brightness of the pixels.
@@
For our high-order maps, designed to have high spatial freedom, we found that a spherical harmonic basis of $\ell_{\mathrm{max}}=4$ and an oversampling factor of 3 (resulting in 62 pixels to sample) produced a good compromise between spatial freedom and computational time. For our low-order maps, designed to compare directly to the null hypothesis of an $n_{\mathrm{max}}=2$ Fourier series, we use a basis of $\ell_{\mathrm{max}}=2$ and an oversampling factor of 3 (corresponding to 16 pixels). The largest-scale eclipse mapping signals may be approximately fitted with an $\ell_{\mathrm{max}}=1$ map, but such a map would not in general be capable of fitting the phase curve outside eclipse. This is because the $n=2$ modes are still significant for the phase curve of a generic map (Cowan & Agol, 2008), so we choose to use the $\ell_{\mathrm{max}}=2$ map to capture these.
We define the pixels for a given map as a vector of brightness values $\mathbf{p}$. The sampled pixel values $\mathbf{p}$ define the vector $\mathbf{c}$ of the coefficients $c_{i}$ of the spherical harmonic representation via a matrix $\mathbf{M}$:
@@
We fit a model $\mathcal{M}$ to the data points $D$ in the light curve using PyMC3 (Salvatier et al., 2016). We sample the posterior distribution of the brightness values of the pixels $\mathbf{p}$ with 250 samples in 8 chains, thinned by a factor of 4. This ensured that the Gelman-Rubin statistic was below 1.1 for our fitted maps. It would be possible to fit the orbital and instrumental parameters at the same time as this map (Hammond et al., 2024), but in this study we assume these parameters are known in advance. The sampling process for one map took around 15 minutes for an $\ell_{\mathrm{max}}=4$ map with 62 pixels using 8 cores on a recent Intel-based HPC server. Calculating the optimal information content of the map with cross-validation requires re-fitting the model around 300 times, resulting in a total runtime of several days. We settled on this number of samples as we did not find a noticeable improvement in mapping performance with more samples.
The $\ell_{\mathrm{max}}=4$ is intended to be the highest-order map that is computationally possible, as we aim for the most degrees of spatial freedom. This is why we do not consider $\ell_{\mathrm{max}}=3$ maps, as we only need a “simplest model” ($\ell_{\mathrm{max}}=2$) to detect a signal, and then a “most complex model” ($\ell_{\mathrm{max}}=4$) to fit the most accurate map. We chose the $\ell_{\mathrm{max}}=4$ map based on the sampling time, as using $\ell_{\mathrm{max}}=5$ maps took several times the sampling time, and sometimes did not sample the posterior fully. Using $\ell_{\mathrm{max}}=5$ maps also produced no noticeable improvement in the optimised map as any small-scale variation due to the $\ell=5$ modes was smoothed out when the information content was optimised.
@@ 2.7 Optimising the spatial scale of fitted maps
We include the information content of the map in the fitting process by penalising the model likelihood $p$ by a factor $2\alpha S$. This provides a functional that balances the model likelihood against the mapped complexity, with the balance set by the free parameter $\alpha$, as originally used in Vogt et al. (1987). We follow the notation in Chen et al. (2024) and formulate the functional as the likelihood penalised by the entropy:
@@ 3.1 Identifying eclipse mapping signals in simulated data
Figure 4 shows the eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ fitted to the $\sigma=$ 150 ppm dataset in Figure 2. The first panel shows the median of the posterior distribution of fitted maps (note that the two-dimensional structure is only constrained on the dayside). We plot the median at every point because we sample the pixels composing the map; this plot therefore shows the median value of the sampled posterior distribution. We confirmed (but do not plot) that this median map is very similar to the maximum likelihood map from the posterior distribution, as discussed in Hammond et al. (2024) This is the case in general for all of the maps in this paper with relatively narrow posterior distributions. This is not the case for high-order maps with very wide posterior distributions, where the maximum likelihood map can be very different to the median map, as it is overfitting to noise and derives a particular set of spurious small-scale features that happen to fit the noise best.
The next two panels of Figure 4 show slices of the posterior distribution of the fitted maps on the dayside through the substellar point, as functions of longitude and latitude. The true map is composed of many high-order spherical harmonic components beyond $\ell=2$, so this fitted $\ell_{\mathrm{max}}=2$ map cannot match the shape of the true map very closely. It does, at least, fit the residual signal accurately, so we suggest that this map is statistically supported, if not especially accurate.
Hammond et al. (2024) considered the ”observable” and “null” spaces of eclipse maps of GCM simulations as defined by Challener & Rauscher (2023), which are the components of the map that do and do not contribute to the mapping signal respectively. We do not consider the observable space in this study as we want to instead highlight the loss of information from fitting a map with a specific set of spherical harmonics, as well as the ability of our optimisation method to access the true spatial scale of the GCM. However, it should be remembered that the black dashed lines showing the “true” GCM maps in Figure 4 and onwards are not exactly accessible by the conversion from light curve to spatial map; see Challener & Rauscher (2023) for more discussion.
@@
The first row of Figure 3 shows the simulated data for $\sigma=$ 150 ppm, with the Fourier series and eclipse map models fitted using an incorrect value for the transit time $t_{0}^{\prime}=t_{0}+10\mathrm{s}$. As we assume a circular orbit, this produces the same offset of $+10\mathrm{s}$ in the assumed eclipse time. This simple change produces a systematic error in the astrophysical model, which means that the eclipse mapping signal and the eclipse map are fitted poorly regardless of the precision of the data. An error in assumed eclipse time could also arise in other ways, such as a slightly eccentric orbit producing an offset in eclipse timing relative to a measured transit time. Fitting the model with this delayed eclipse timing results in an eclipse ingress that is too high, and an eclipse egress that is too low. The model wrongly adjusts the eclipse map to compensate for this error. Williams et al. (2006) showed how a timing error can produce a spurious longitudinal offset in a map from an isolated eclipse. We suggest that a longitudinal offset is not introduced in this case because the low-order longitudinal structure is strongly constrained by the out-of-eclipse phase curve (which is not significantly affected by the small timing error). The model therefore adjust the latitudinal structure instead to attempt to better fit the residual signal in the eclipse shape.
@@ 3.3 Optimising the information content of fitted maps
Having used a simple eclipse mapping model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ to establish the presence of an eclipse mapping signal in the dataset with $\sigma=$ 150 ppm in Figure 2, we can now fit a map with more complex structure. Figure 7 shows an eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=4)$ fitted to this dataset using spherical harmonics up to order $\ell_{\mathrm{max}}=4$. This high-order map is able to fit the true map in theory, but in practice it overfits a wide range of solutions, resulting in an impractically large uncertainty. The large uncertainty on the fitted position of the brightest point, at $(56^{+65}_{-58})^{\circ}$, reflects the range of different small-scale peaks fitted with the high-order harmonics, and shows the limitation of using this metric for anything but the largest-scale maps.
@@
Figure 9: The thermal emission from 5 to 10.5 $\mu$m from the hot Jupiter WASP-43b observed by JWST MIRI/LRS (Bell et al., 2024). The layout of panels is the same as Figure 2. The black points show the first eclipse and the green points show the second eclipse, binned every 10 points. The eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=2)$ fits the eclipse shape much better than the Fourier series model $\mathcal{M}(n_{\mathrm{max}}=2)$. This is confirmed by the higher cross-validation score for the eclipse map model for this dataset shown in Figure 12 and listed in Table 2. The data points in the right-hand panel appear slightly different to the data points in Hammond et al. (2024) because here we bin them with a slightly longer time cadence.